11 11 TEX L^{A}TEX

The Linux Thai HOWTO

Poonlap Veeratanabutr, poon-v@fedu.uec.ac.jp v0.4, 4 August 1998

This document describes how to use Thai language with Linux. This will cover setting Thai fonts, Thai keyboard and some Thai applications.

Contents

1 Introduction 1

2 Thai Input and Output 2

2.1 Linux console . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.1.1 Thai fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.1.2 Keyboard layout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

2.2 X Window system . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.2.1 Thai fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.2.2 Thai keyboard layout There are two ways to map Thai keyboard on X window, using

X Keyboard Extension (XKB) and using xmodmap. Please select how you map Thai keyboard. I recommend using XKB. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

3 Applications with Thai language 13

3.1 Some X applications and Thai language . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 3.1.1 txterm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 3.1.2 Emacs, Mule . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 3.1.3 vi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 3.1.4 pine . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.1.5 Netscape . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.1.6 Ss, Simple thai word Separator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.1.7 Xzthai, X keyboard configurator + simple editor . . . . . . . . . . . . . . . . . . . . . 15

3.2 Printing Thai document . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.2.1 Latex and Thai language . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.2.2 Latex’s configuration for Thai language . . . . . . . . . . . . . . . . . . . . . . . . . . 16 3.2.3 Use Thai LaTeX filter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 3.2.4 Editing LaTex source file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

3.3 X Application Resources . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 3.4 Thai Extension for Linux (TE) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

4 References and FTP sites 18 4.1 Other documents of relevance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 4.2 Thai related stuffs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 4.3 FTP and Web sites . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

5 Acknowledgments and Copyright 20

1 Introduction

It’s about one year that I didn’t update this document. There were a lot of movement in Thai computing and using Linux in Thailand. For example, Linux boxes are used as server in many schools in Thailand

<http://www.school.net.th>.

The purpose of this document is to show how to set your Linux to use Thai language. I use Linux RedHat 5.0 as I wrote this document, so directories which I mention in this document may be different from other distribution. First I would like to talk about Thai standard character set. Thai standard character set is TIS-620. There are also other Thai standard character sets such as ISO-IR-166, CP874, etc. Please see http://www.inet.co.th/cyberclub/trin/thairef/ <http://www.inet.co.th/cyberclub/trin/thairef/> for further information about Thai standard character set. TIS-620 is 8-bit character set. It has the same range as ISO-8859-1, so we can use apllications that support ISO-8859-1 character set also. Although we can use Thai language with apllications that support ISO-8859-1 character set, but it does not mean those applications support Thai language.

Thai characters are different from English characters. There is a variation of position, normal position, character can be on other character, character can be under other character. There is no space between words. These are some problems in developing Thai supported application.

You can find the lastest version of Thai-HOWTO document from http://www.fedu.uec.ac.jp/ZzzThai/Linux

<http://www.fedu.uec.ac.jp/ZzzThai/Linux>. Your comment is welcome.

2 Thai Input and Output

2.1 Linux console

Thai characters do not display properly in Linux console. If you mainly use X window, you may pass this section.

2.1.1 Thai fonts

You can obtain Linux Thai console fonts which created by Mr. Phaisarn Techajaruwong from ZzzThai ftp

site <ftp://ftp.fedu.uec.ac.jp/pub/thai/UEC/ZzzThai/Sofware/Linux/fonts/console>

For example, there is a font name "phaisarn.psf". Put it in /usr/lib/kbd/consolefonts/ directory. Then, you can load the new font from Linux console by command

     %setfont phaisarn.psf

2.1.2 Keyboard layout

You can set keyboard behavior as you like by using loadkeys command. Usually, you use loadkeys to load the file located in /usr/lib/kbd/keytables. You can create a US/Thai keyboard-map file and save it in this directory. Here is a sample.

     keycode   0 =
     keycode   1 = Escape           Escape
             alt     keycode   1 = Meta_Escape
     keycode   2 = +one              exclam          +0x0e5          plus
             alt     keycode   2 = Meta_one
             alt shift keycode 2 = Meta_exclam
     keycode   3 = +two              at              +slash          0x0f1
             control keycode   3 = nul
             control shift keycode 3 = nul
             alt     keycode   3 = Meta_two
             alt shift keycode 3 = Meta_at
     keycode   4 = +three            numbersign      +underscore     0x0f2
             control keycode   4 = Escape
             alt     keycode   4 = Meta_three
             alt shift keycode 4 = Meta_numbersign
     keycode   5 = +four             dollar          +0x0c0          0x0f3

2. Thai Input and Output 3

control keycode   5 = Control_backslash
alt     keycode   5 = Meta_four
alt shift keycode 5 = Meta_dollar
keycode   6 = +five             percent         +0x0b6          0x0f4
control keycode   6 = Control_bracketright
alt     keycode   6 = Meta_five
alt shift keycode 6 = Meta_percent
keycode   7 = +six              asciicircum     +0x0d8          0x0d9
control keycode   7 = Control_asciicircum
alt     keycode   7 = Meta_six
alt shift keycode 7 = Meta_asciicircum
keycode   8 = +seven            ampersand       +0x0d6          0x0df
control keycode   8 = Control_underscore
alt     keycode   8 = Meta_seven
keycode   9 = +eight            asterisk        +0x0a4          0x0f5
control keycode   9 = Delete
alt     keycode   9 = Meta_eight
keycode  10 = +nine             parenleft       +0x0b5          0x0d6
alt     keycode  10 = Meta_nine
keycode  11 = +zero             parenright      +0x0a8          0x0f7
alt     keycode  11 = Meta_zero
keycode  12 = +minus            underscore      +0x0a2          0x0f8
control keycode  12 = Control_underscore
control shift keycode 12 = Control_underscore
alt     keycode  12 = Meta_minus
keycode  13 = +equal            plus            +0x0aa          0x0f9
alt     keycode  13 = Meta_equal
keycode  14 = Delete           Delete           Delete          Delete
alt     keycode  14 = Meta_Delete
keycode  15 = Tab              Tab              Tab             Tab
alt     keycode  15 = Meta_Tab
keycode  16 = +q        Q       +0x0e6          0x0f0
keycode  17 = +w        W       +0x0e4          quotedbl
keycode  18 = +e        E       +0x0d3          0x0ae
keycode  19 = +r        R       +0x0be          0x0b1
keycode  20 = +t        T       +0x0d0          0x0b8
keycode  21 = +y        Y       +0x0d1          0x0ed
keycode  22 = +u        U       +0x0d5          0x0ea
keycode  23 = +i        I       +0x0c3          0x0b3
keycode  24 = +o        O       +0x0b9          0x0cf
keycode  25 = +p        P       +0x0c2          0x0ad
keycode  26 = +bracketleft      braceleft       +0x0ba          0x0b0
control keycode  26 = Escape
alt     keycode  26 = Meta_bracketleft
alt shift keycode 26 = Meta_braceleft
keycode  27 = +bracketright     braceright      +0x0c5          comma
control keycode  27 = Control_bracketright
alt     keycode  27 = Meta_bracketright
alt shift keycode 27 = Meta_braceright
keycode  28 = Return            Return          Return          Return
alt     keycode  28 = 0x080d

keycode  29 = Control           Control         Control         Control
keycode  30 = +a        A       +0x0bf          0x0c4
keycode  31 = +s        S       +0x0cb          0x0a6
keycode  32 = +d        D       +0x0a1          0x0af
keycode  33 = +f        F       +0x0b4          0x0e2
keycode  34 = +g        G       +0x0e0          0x0ac
keycode  35 = +h        H       +0x0e9          0x0e7
keycode  36 = +j        J       +0x0e8          0x0eb
keycode  37 = +k        K       +0x0d2          0x0c9
keycode  38 = +l        L       +0x0ca          0x0c8
keycode  39 = +semicolon        colon           +0x0c7          0x0ab
        alt     keycode  39 = Meta_semicolon
keycode  40 = +apostrophe       quotedbl        +0x0a7          period
        control keycode  40 = Control_g
        alt     keycode  40 = Meta_apostrophe
keycode  41 = +grave            asciitilde      +minus          percent
        control keycode  41 = nul
        alt     keycode  41 = Meta_grave
keycode  42 = Shift            Shift            Shift           Shift
keycode  43 = +backslash        bar             +0x0a3          0x0a5
        control keycode  43 = Control_backslash
        alt     keycode  43 = Meta_backslash
        alt shift keycode 43 = Meta_bar
keycode  44 = +z        Z       +0x0bc          parenleft
keycode  45 = +x        X       +0x0bb          parenright
keycode  46 = +c        C       +0x0e1          0x0a9
keycode  47 = +v        V       +0x0cd          0x0ce
keycode  48 = +b        B       +0x0d4          0x0da
keycode  49 = +n        N       +0x0d7          0x0ec
keycode  50 = +m        M       +0x0b7          question
keycode  51 = +comma    less    +0x0c1          0x0b2
        alt     keycode  51 = Meta_comma
        alt shift keycode 51 = Meta_less
keycode  52 = +period           greater         +0x0e3          0x0cc
        alt     keycode  52 = Meta_period
        alt shift keycode 52 = Meta_greater
keycode  53 = +slash            question        +0x0bd          0x0c6
        control keycode  53 = Delete
        alt     keycode  53 = Meta_slash
keycode  54 = Shift            Shift            Shift           Shift
keycode  55 = KP_Multiply
keycode  56 = Alt              Alt              Alt             Alt
keycode  57 = space            space            space           space
        control keycode  57 = nul
        alt     keycode  57 = Meta_space
keycode  58 = Caps_Lock        Caps_Lock        Caps_Lock       Caps_Lock
keycode  59 = F1               F11              Console_13
        control keycode  59 = F1
        alt     keycode  59 = Console_1
        control alt     keycode  59 = Console_1
keycode  60 = F2               F12              Console_14

control keycode  60 = F2
alt     keycode  60 = Console_2
control alt     keycode  60 = Console_2
keycode  61 = F3               F13              Console_15
control keycode  61 = F3
alt     keycode  61 = Console_3
control alt     keycode  61 = Console_3
keycode  62 = F4               F14              Console_16
control keycode  62 = F4
alt     keycode  62 = Console_4
control alt     keycode  62 = Console_4
keycode  63 = F5               F15              Console_17
control keycode  63 = F5
alt     keycode  63 = Console_5
control alt     keycode  63 = Console_5
keycode  64 = F6               F16              Console_18
control keycode  64 = F6
alt     keycode  64 = Console_6
control alt     keycode  64 = Console_6
keycode  65 = F7               F17              Console_19
control keycode  65 = F7
alt     keycode  65 = Console_7
control alt     keycode  65 = Console_7
keycode  66 = F8               F18              Console_20
control keycode  66 = F8
alt     keycode  66 = Console_8
control alt     keycode  66 = Console_8
keycode  67 = F9               F19              Console_21
control keycode  67 = F9
alt     keycode  67 = Console_9
control alt     keycode  67 = Console_9
keycode  68 = F10              F20              Console_22
control keycode  68 = F10
alt     keycode  68 = Console_10
control alt     keycode  68 = Console_10
keycode  69 = Num_Lock
keycode  70 = Scroll_Lock      Show_Memory      Show_Registers
control keycode  70 = Show_State
alt     keycode  70 = Scroll_Lock
keycode  71 = KP_7
alt     keycode  71 = Ascii_7
keycode  72 = KP_8
alt     keycode  72 = Ascii_8
keycode  73 = KP_9
alt     keycode  73 = Ascii_9
keycode  74 = KP_Subtract
keycode  75 = KP_4
alt     keycode  75 = Ascii_4
keycode  76 = KP_5
alt     keycode  76 = Ascii_5
keycode  77 = KP_6

        alt     keycode  77 = Ascii_6
keycode  78 = KP_Add
keycode  79 = KP_1
        alt     keycode  79 = Ascii_1
keycode  80 = KP_2
        alt     keycode  80 = Ascii_2
keycode  81 = KP_3
        alt     keycode  81 = Ascii_3
keycode  82 = KP_0
        alt     keycode  82 = Ascii_0
keycode  83 = KP_Period
        altgr   control keycode  83 = Boot
        control alt     keycode  83 = Boot
keycode  84 = Last_Console
keycode  85 =
keycode  86 = less             greater          bar
        alt     keycode  86 = Meta_less
keycode  87 = F11              F11              Console_23
        control keycode  87 = F11
        alt     keycode  87 = Console_11
        control alt     keycode  87 = Console_11
keycode  88 = F12              F12              Console_24
        control keycode  88 = F12
        alt     keycode  88 = Console_12
        control alt     keycode  88 = Console_12
keycode  89 =
keycode  90 =
keycode  91 =
keycode  92 =
keycode  93 =
keycode  94 =
keycode  95 =
keycode  96 = KP_Enter
keycode  97 = Control
keycode  98 = KP_Divide
keycode  99 = Control_backslash
        control keycode  99 = Control_backslash
        alt     keycode  99 = Control_backslash
keycode 100 = AltGr_Lock
keycode 101 = Break
keycode 102 = Find
keycode 103 = Up
keycode 104 = Prior
        shift   keycode 104 = Scroll_Backward
keycode 105 = Left
        alt     keycode 105 = Decr_Console
keycode 106 = Right
        alt     keycode 106 = Incr_Console
keycode 107 = Select
keycode 108 = Down
keycode 109 = Next

        shift   keycode 109 = Scroll_Forward
keycode 110 = Insert
keycode 111 = Remove
        altgr   control keycode 111 = Boot
        control alt     keycode 111 = Boot
keycode 112 =
keycode 113 =
keycode 114 =
keycode 115 =
keycode 116 =
keycode 117 =
keycode 118 =
keycode 119 =
keycode 120 =
keycode 121 =
keycode 122 =
keycode 123 =
keycode 124 =
keycode 125 =
keycode 126 =
keycode 127 =
string F1 = "\033[[A"
string F2 = "\033[[B"
string F3 = "\033[[C"
string F4 = "\033[[D"
string F5 = "\033[[E"
string F6 = "\033[17~"
string F7 = "\033[18~"
string F8 = "\033[19~"
string F9 = "\033[20~"
string F10 = "\033[21~"
string F11 = "\033[23~"
string F12 = "\033[24~"
string F13 = "\033[25~"
string F14 = "\033[26~"
string F15 = "\033[28~"
string F16 = "\033[29~"
string F17 = "\033[31~"
string F18 = "\033[32~"
string F19 = "\033[33~"
string F20 = "\033[34~"
string Find = "\033[1~"
string Insert = "\033[2~"
string Remove = "\033[3~"
string Select = "\033[4~"
string Prior = "\033[5~"
string Next = "\033[6~"
string F21 = ""
string F22 = ""
string F23 = ""
string F24 = ""

     string F25 = ""
     string F26 = ""

Suppose you save this file as thai.map. From Linux console, use command loadkeys to load thai.map.

     %loadkeys thai.map

You can switch to Thai keyboard by pressing the right Alt key. If you want to switch the keyboard back, press the right Alt key again.

2.2 X Window system

2.2.1 Thai fonts

You can obtain Thai fonts in bdf format or pcf format from internet. You can also use scalable fonts such as Type1 or TrueType fonts. But I will not describe about these.

Installing Thai fonts You must log in as root. Let’s put Thai fonts in /usr/X11R6/lib/X11/fonts/misc/, this is a default font path. Change directory to

/usr/X11R6/lib/X11/fonts/misc/ and run command
     %mkfontdir
     %xset fp rehash

If you put Thai fonts in different directory, you must use xset to add the new font path. Please see man-page for further information. You can check the new fonts by running command xlsfonts and see whether there are Thai fonts or not. If you can not see any Thai fonts from this command, you may need to restart X window.

2.2.2 Thai keyboard layout There are two ways to map Thai keyboard on X window, using X Keyboard Extension (XKB) and using xmodmap. Please select how you map Thai keyboard. I recommend using XKB.

XKB and Thai keyboard layout. Beginning with XFree86 3.1.2D, you can use the new X11R6.1 XKEY-BOARD extension to manage the keyboard layout. This is very helpful.

During X server configuration with xf86config you will be asked about XKB, if you want to to set Thai keyboard layout for your system, say yes. There are a list of pre-configured keymaps. Choose Standard 101-key, Thai encoding.

XF86Setup is the graphical X server configuration utility for XFree86 X server. It is easier than traditional xf86config. You can select a keyboard layout easily with this tool.

Ther are many choices of keyboard switch key to select. The default is LeftAlt+RightShift switch to Thai and LeftAlt+LeftShift switch to US. You can type Thai characters in any applications which support ISO-8859-1 character set, but don’t forget to use Thai fonts with those applications too.

I found that pre-configured keymaps that came with XFree86-3.2 is not correct. You may not be able to type THO THUNG which located at " 5 key " . To fix this problem, you should add the line

         key <AE05> {        [], [   paragraph,      ocircumflex     ]       };
in the file /usr/X11R6/lib/X11/xkb/symbols/th as the example.
     .................
         key <AE03> {        [], [       minus,      ograve          ]       };
         key <AE04> {        [], [      Agrave,      oacute          ]       };
         key <AE05> {        [], [   paragraph,      ocircumflex     ]       };
         key <AE06> {        [], [    Ooblique,      Ugrave          ]       };
         key <AE07> {        [], [  Odiaeresis,      ssharp          ]       };
     .................

You can not type SORUSI also. Please change the line from

         key <AC08> {        [], [      Ograve,      eacute          ]       };

to

         key <AC08> {        [], [      Ograve,      Eacute          ]       };

Note that eacute is equal to MAITHO and Eacute is equal to SORUSI.

There are also XKB extension utilities such as setxkbmap, xkbcomp, etc. Please see man-page for more information. I recommend to use xkbvled. The leds will be on when you are using Thai keyboard so you can know your keyboard’s status.

The following is part of XF86Config file about keyboard section. If you want to configure the keyboard by hand, change the content of /usr/X11R6/lib/X11/XF86Config as an example below. This configuration uses the default toggle key.

     Section "Keyboard"
        Protocol        "Standard"
        AutoRepeat      500 5
        LeftAlt         Meta
        RightAlt        Meta
        ScrollLock      Compose
        RightCtl        Control
     #  XkbDisable
        XkbKeycodes     "xfree86"
        XkbTypes        "default"
        XkbCompat       "default"
        XkbSymbols      "us(pc101)"
        XkbGeometry     "pc"
        XkbRules        "xfree86"
        XkbModel        "pc101"
        XkbLayout       "th"
     EndSection

If you use XKB extension, Thai keyboard mapping with xmodmap may not work. See XF86Config man-page for mor information.

Thai keyboard layout with xmodmap You can use the utility xmodmap to map Thai keyboard. Normally xmodmap is used to load a keyboard configured file. For most Linux distributions, when you start X window with startx, X server will find .Xmodmap in /usr/X11R6/lib/X11/xinit/ first. If .Xmodmap does not exist, X server will find .Xmodmap in your home directory. Please see the content of

/usr/X11R6/lib/X11/xinit/xinitrc.

The following is the sample of .Xmodmap for Thai Kedmanee keyboard layout.

!
! Linux/XFree86 Thai Kedmanee layout (based on US keyboard)
! Use ScrollLock to switch to Thai keyboard.
! This file will work with XFree86 only.
!
keycode 0x09 =  Escape
keycode 0x43 =  F1
keycode 0x44 =  F2
keycode 0x45 =  F3
keycode 0x46 =  F4
keycode 0x47 =  F5
keycode 0x48 =  F6
keycode 0x49 =  F7
keycode 0x4A =  F8
keycode 0x4B =  F9
keycode 0x4C =  F10
keycode 0x5F =  F11
keycode 0x60 =  F12
keycode 0x6F =  Print
keycode 0x4E =  Mode_switch     XF86ModeLock
keycode 0x6E =  Pause
keycode 0x31 =  grave           asciitilde      minus           percent
keycode 0x0A =  1               exclam          0x0e5           plus
keycode 0x0B =  2               at              slash           0x0f1
keycode 0x0C =  3               numbersign      underscore      0x0f2
keycode 0x0D =  4               dollar          0x0c0           0x0f3
keycode 0x0E =  5               percent         0x0b6           0x0f4
keycode 0x0F =  6               asciicircum     0x0d8           0x0d9
keycode 0x10 =  7               ampersand       0x0d6           0x0df
keycode 0x11 =  8               asterisk        0x0a4           0x0f5
keycode 0x12 =  9               parenleft       0x0b5           0x0f6
keycode 0x13 =  0               parenright      0x0a8           0x0f7
keycode 0x14 =  minus           underscore      0x0a2           0x0f8
keycode 0x15 =  equal           plus            0x0aa           0x0f9
keycode 0x33 =  backslash       bar             0x0a3           0x0a5
keycode 0x16 =  BackSpace
keycode 0x6A =  Insert
keycode 0x61 =  Home
keycode 0x63 =  Prior
keycode 0x4D =  Num_Lock
keycode 0x70 =  KP_Divide
keycode 0x3F =  KP_Multiply

keycode 0x52 =  KP_Subtract
keycode 0x17 =  Tab
keycode 0x18 =  q               Q               0x0e6           0x0f0
keycode 0x19 =  w               W               0x0e4           quotedbl
keycode 0x1A =  e               E               0x0d3           0x0ae
keycode 0x1B =  r               R               0x0be           0x0b1
keycode 0x1C =  t               T               0x0d0           0x0b8
keycode 0x1D =  y               Y               0x0d1           0x0ed
keycode 0x1E =  u               U               0x0d5           0x0ea
keycode 0x1F =  i               I               0x0c3           0x0b3
keycode 0x20 =  o               O               0x0b9           0x0cf
keycode 0x21 =  p               P               0x0c2           0x0ad
keycode 0x22 =  bracketleft     braceleft       0x0ba           0x0b0
keycode 0x23 =  bracketright    braceright      0x0c5           comma
keycode 0x24 =  Return
keycode 0x6B =  Delete
keycode 0x67 =  End
keycode 0x69 =  Next
keycode 0x4F =  KP_7
keycode 0x50 =  KP_8
keycode 0x51 =  KP_9
keycode 0x56 =  KP_Add
keycode 0x42 =  Caps_Lock
keycode 0x26 =  a               A               0x0bf           0x0c4
keycode 0x27 =  s               S               0x0cb           0c0a6
keycode 0x28 =  d               D               0x0a1           0x0af
keycode 0x29 =  f               F               0x0b4           0x0e2
keycode 0x2A =  g               G               0x0e0           0x0ac
keycode 0x2B =  h               H               0x0e9           0x0e7
keycode 0x2C =  j               J               0x0e8           0x0eb
keycode 0x2D =  k               K               0x0d2           0x0c9
keycode 0x2E =  l               L               0x0ca           0x0c8
keycode 0x2F =  semicolon       colon           0x0c7           0x0ab
keycode 0x30 =  apostrophe      quotedbl        0x0a7           period
keycode 0x53 =  KP_4
keycode 0x54 =  KP_5
keycode 0x55 =  KP_6
keycode 0x32 =  Shift_L
keycode 0x34 =  z               Z               0x0bc           parenleft
keycode 0x35 =  x               X               0x0bb           parenright
keycode 0x36 =  c               C               0x0e1           0x0a9
keycode 0x37 =  v               V               0x0cd           0x0ce
keycode 0x38 =  b               B               0x0d4           0x0da
keycode 0x39 =  n               N               0x0d7           0x0ec
keycode 0x3A =  m               M               0x0b7           question
keycode 0x3B =  comma           less            0x0c1           0x0b2
keycode 0x3C =  period          greater         0x0e3           0x0cc
keycode 0x3D =  slash           question        0x0bd           0x0c6
keycode 0x3E =  Shift_R
keycode 0x62 =  Up
keycode 0x57 =  KP_1

     keycode 0x58 =  KP_2
     keycode 0x59 =  KP_3
     keycode 0x6C =  KP_Enter
     keycode 0x25 =  Control_L
     keycode 0x40 =  Alt_L           Meta_L
     keycode 0x41 =  space
     keycode 0x71 =  Alt_R           Meta_R
     keycode 0x6D =  Control_R
     keycode 0x64 =  Left
     keycode 0x68 =  Down
     keycode 0x66 =  Right
     keycode 0x5A =  KP_0
     keycode 0x5B =  KP_Decimal
     clear Shift
     clear Lock
     clear Control
     clear Mod1
     clear Mod2
     clear Mod3
     clear Mod4
     clear Mod5
     add    Shift   = Shift_L Shift_R
     add    Lock    = Caps_Lock
     add    Control = Control_L Control_R
     add    Mod1    = Alt_L Alt_R
     add    Mod2    = Mode_switch

Just put .Xmodmap in your home directory will be OK. When you start X window, X server will load this file.

You can also load .Xmodmap from command line.

     %xmodmap ~/.Xmodmap

In above .Xmodmap file, US/Thai switch key is assigned to keycode 0x4E (78), Scroll Lock key, with the statement

     keycode 0x4E =  Mode_switch     XF86ModeLock

XF86ModeLock is the special keysym for XFree86 X server. If you don’t add this keysym, you have to hold the scroll lock key while you are typing Thai characters. Note that if you use commercial X server, some keycodes are different. You may have to map keyboard by yourself. See man-pages of X and xev for further information.

Note: If you are using XFree86 version 3.1.2D or later, you need to add the line XkbDisable in keyboard section of XF86Config file. You may config the keyboard section like the following sample.

     Section "Keyboard"
        Protocol        "Standard"
        AutoRepeat      500 5

        LeftAlt         Meta
        RightAlt        ModeShift
        ScrollLock      ModeLock
        RightCtl        Compose
        XkbDisable
     EndSection

3 Applications with Thai language

This is the tricky part. Most applications support ISO-8859-1 character set. For example, emacs can display ISO-8859-1 character. If we set emacs to display ISO-8859-1 and use Thai font, you can edit Thai document with emacs. But this is not a good policy. You should avoid using this trick as possible. What we need is Thai locale or Thai supported applications to manage these things.

To make X window application displays Thai font, you should run the application with -fn option. For example,

     %xedit -fn thai8x16

Note that thai8x16 is just a one of Thai font names. You can see all available fonts by command xlfonts. If you don’t want to fill -fn option every time you run application. You should set Thai font in your

˜ /.Xdefaults or ˜ /.Xresources like this
     XTerm*font:     thai8x16

3.1 Some X applications and Thai language

3.1.1 txterm

txterm is Thai version of xterm. There are several programs running under xterm such as shell, pine, vi, less, etc. We can type Thai characters without any problems with txterm. Txterm also provides its own Thai input method by pressing " F1 " key. Txterm will use fonts thai9x13 as default Thai font. You can change this by add -fn option.

You can get txterm from Thaigate <http://thaigate.nacsis.ac.jp> or ZzzThai <ftp://ftp.fedu.uec.
ac.jp/pub/thai/UEC/ZzzThai/Software/Linux>.
bash shell :

Normally, shell accepts only ASCII character set. To type Thai characters in shell command line, you should set environment LC_CTYPE to iso_8859_1. I don’t set LC_CTYPE environment variable to iso_8859_1 because this environment variable will effect other applications too. With bash shell, you can specify which environment variable to be passed to the application. For example, I can make a fake Thai X terminal with this syntax.

          LC_CTYPE=iso_8859_1 xterm -fn thai8x16

This xterm display Thai characters well, but not good for typing Thai characters. I strongly recommend you to use txterm.

ls :

If you name a filename in Thai. Issue the command as

          ls -N

You may set alias in ˜ /.bashrc or ˜ /.cshrc, so you can type ls without option. If you don’t use ls with -N option, you may see Thai filename as ?????.

3.1.2 Emacs, Mule

Mule stands for " Multilingual Enhancement to GNU Emacs " . It has the same functions as emacs and supports many languages. Mule provides its own input method, so you don’t need any configuration for typing Thai. You needs only Thai fonts for mule which you can get from, ZzzThai <ftp://ftp.fedu.uec.ac.

jp/pub/thai/UEC/ZzzThai/Software/Linux/fonts/> or Etl site <ftp://etlport.etl.go.jp/pub/mule/

fonts/ETL/>. These Thai fonts are fixed width fonts.

You need some configuration for mule. Puts the following lines in your .emacs.

     ;;
     ;; Thai System, add in .emacs
     ;;
     (set-file-coding-system-for-read '*tis620*)
     (set-default-file-coding-system '*tis620*)
     (set-display-coding-system '*tis620*)
     (set-keyboard-coding-system '*tis620*)
     (setq-deafault quail-current-package (assoc "thai" quail-package-alist))

Add the following lines in .Xdefaults.

     !
     ! Emacs, Mule - Font menu
     !
     Emacs*FontSetList: thai14, thai16, thai24
     Emacs*FontSet-thai14:\
             -etl-fixed-medium-r-normal--14-140-72-72-m-70-tis620.2529-1
     Emacs*FontSet-thai16:\
             -etl-fixed-medium-r-normal--16-160-72-72-m-80-tis620.2529-1
     Emacs*FontSet-thai24:\
             -etl-fixed-medium-r-normal--24-500-72-72-m-120-tis620.2529-1

When you hold shift key and press left mouse’s button, you can select Thai fonts to display in mule window. To type Thai characters, press " Ctrl + ] " . To type English, press " Ctrl + ] " again.

You can get mule from ElectroTechnical Laboratory(ETL) <ftp://etlport.etl.go.jp/pub/mule/>

3.1.3 vi

Vi should be run on txterm.

3.1.4 pine

In the past, we could not send 8-bit characters through E-mail. Now, although mail transfer agent can handle 8-bit characters but some old mail transfer agent can not. We can send Thai e-mail by using e-mail application that supports MIME (Multipurpose Internet Mail Extensions) E-mail applications that support MIME are pine, elm, Netscape mail, etc.

Put the following definition in your ˜ /.pinerc file:

     character-set=ISO-8859-1

This can also be set via the Setup option in pine window. You can find it under Config. You can read Thai news from pine, too.

Pine should be run in txterm.

3.1.5 Netscape

If you have Thai fonts in your system. Just set Thai font from preference. Thai fonts will appear in User defined area. See http://www.fedu.uec.ac.jp/ZzzThai/unix <http://www.fedu.uec.ac.jp/ZzzThai/unix> for setting Thai language on Netscape.

Some movement about Thai Mozilla project at http://members.xoom.com/inThai/mozthai.html <http://

members.xoom.com/inThai/mozthai.html>.

3.1.6 Ss, Simple thai word Separator

ss is a dictionary based Thai word separation program similar to cttex. It can be used to insert a configurable string between Thai words. It can also show words that cannot be found in the dictionary. More words can be added to the dictionary. Developed by Mr.Teera Kittichareonpot.

We can use this program to insert < WBR > tag between Thai words in html file. Browser will display Thai homepage better than normal html document.

3.1.7 Xzthai, X keyboard configurator + simple editor

Xzthai, this is the Tcl/Tk application for mapping Thai keyboard on any keyboard with graphical user interface. Also provides simple editor and keyboard layout figure. It actually uses xmodmap program in background to map Thai keyboard. This may be useful for commercial X server and X server on UNIX.

3.2 Printing Thai document

Thai2ps is used to convert plain text file to postscript file.You can use ghostscript(gs) to print your Thai document. For better quality document, you have to use (La)TeX.

3.2.1 Latex and Thai language

Dr. Manop Wongsaisuwan first tried to use Thai language with latex. He wrote some perl script as filter for latex source code that contains Thai language. Then pass the result to latex. Mr. Vuthichai Ampornaramvech used this concept and wrote a program in C language, cttex, to handle this. It runs faster and makes Thai word segmentation based on dictionary. Cttex also fixes the position of Thai characters in word, so SARA and WANNAYUK will be placed in the beautiful position.

You can find Thai latex filter from http://thaigate.nacsis.ac.jp/files/ttex.html <http://thaigate.nacsis.

ac.jp/files/ttex.html>.

3.2.2 Latex’s configuration for Thai language

You must have latex installed in your computer. First, download Thai postscript (Type1) fonts, tfm fonts and Thai style file. These fonts are needed by Latex. This is the list of files you should download.

     tfm fonts:
         dbtt.tfm    dbttb.tfm   dbttbi.tfm  dbtti.tfm
     postscript fonts:
         dbtt.pfa    dbttb.pfa   dbttbi.pfa  dbtti.pfa
     style files:
         thai.sty sakka.sty
     Thai Latex filter:
         cttex
     Sample Latex file:
         ttex.ttex test.ttex
There is latex’s directory at /usr/lib/texmf/texmf/ (RedHat 5.0). I will call /usr/lib/texmf/texmf/

as " $texroot " . We will concentrate at $texroot/texmf/ directory. In $texroot/texmf/ directory, there are many files about tex’s configuration. You have to edit files in dvips, fonts, tex subdirectories.

Add the following lines to $texroot/texmf/dvips/misc/psfonts.map
     dbtt  DBThaiText <dbtt.pfa
     dbttb DBThaiTextBold <dbttb.pfa
     dbttbi DBThaiTextBoldItalic <dbttbi.pfa
     dbtti DBThaiTextItalic <dbtti.pfa

Make a new directories and copy files to the appropriate directories.

     %mkdir /usr/lib/texmf/texmf/fonts/tfm/public/thai
     %mkdir /usr/lib/texmf/texmf/fonts/type1/public/thai
     %mkdir /usr/lib/texmf/texmf/tex/generic/thai
     %cp *.tfm /usr/lib/texmf/texmf/fonts/tfm/public/thai
     %cp *.pfa /usr/lib/texmf/texmf/fonts/type1/public/thai
     %cp *.sty /usr/lib/texmf/texmf/tex/generic/thai

Run texhash or MakeTeXls-R(in some system) to update Tex database.

     %/usr/bin/texhash
     texhash: updating /usr/lib/texmf/texmf/ls-R ...
     texhash: Done.

3.2.3 Use Thai LaTeX filter

We can use cttex as filter like this,

     %cttex < ttex.ttex > ttex.tex
     C-TTeX $Revivsion: 1.15 $
     Usage : cttex [cutcode] < infile > outfile
     Usage : cutcode=0 forces operation in HTML mode.
     Build-in dictionary size: 9945 words
      343
     Done
     %latex ttex.tex
     ...
     %xdvi ttex.dvi

You can convert dvi file to postscript file by,

     %dvips -o ttex.ps ttex.dvi

Finally, you can print ttex.ps by using gs or lpr. You must configure printer before printing. See man-pages of printcap, gs, lpr for more information.

3.2.4 Editing LaTex source file

For new latex user, lyx is helpful. But I recommend to use mule to edit Thai latex source file because mule supports Thai language and it is a powerful editor. You may take a look a Thai Latex tutorial

<http://www.fedu.uec.ac.jp/ZzzThai/Linux/ttex>.

3.3 X Application Resources

Because Xt based applications allow user to configure the applications by resources. We can make the menu or label to be Thai language.

For example, if you want xman to display Thai labels. You may add these lines in your .Xdefaults

     ......
     !!  Xman section
     Xman*Font:                          thai8x16
     Xman*helpButton.Label:              ^{a}èÇÂ
     Xman*quitButton.Label:              ÍÍ¡
     Xman*manpageButton.Label:           ÙèÁ×^{a}
     ......

You can use the same idea to set window manager to be more Thai environment too.

3.4 Thai Extension for Linux (TE)

Thai Extension for Linux is a installation package comes with applicaions and Thai fonts. You don’t have to configure Linux system and applications by yourself. Let TE do config-

uration task for you.  After installation, you can use Thai language suddenly.  Get TE

from ftp://fedu.uec.ac.jp/pub/thai/UEC/ZzzThai/Software/Linux/ <ftp://fedu.uec.ac.jp/pub/thai/

UEC/ZzzThai/Software/Linux/>

4 References and FTP sites

4.1 Other documents of relevance

The HOWTOs ought to be available from all mirrors of sunsite.unc.edu. The Linux Danish/International HOWTO by Niels Kristian Bech Jensen The Linux Cyrillic HOWTO by Alexander L. Belikoff The Keystroke mini-HOWTO by Zenon Fortuna. The Locales mini-HOWTO by Peeter Joot. (This one is mainly for developers.) The ISO-8859-1 FAQ and Programming for Internationalization FAQ (plus much more) by Michael Gschwind

is available from his homepage <http://www.vlsivie.tuwien.ac.at/mike/i18n.html>.

4.2 Thai related stuffs

" NACSIS R& D Thai Project Page " ,http://thaigate.nacsis.ac.jp <http://thaigate.nacsis.ac.jp> • Information about Thai computing.

• Discussion groups in Thai language, such as thai-l (Thai Mailing list), Thai news, etc.

• Thai references and Thai softwares.

• Thai Latex filter. " ZzzThai Project " , http://www.fedu.uec.ac.jp <http://www.fedu.uec.ac.jp/ZzzThai/>

• Most softwares and Thai fonts introduced here can download from ZzzThai.

• Describe how to use Thai with 3 main computer platforms, UNIX like, Windows and Mac.

• Linux information at http://www.fedu.uec.ac.jp/ZzzThai/Linux <http://www.fedu.uec.ac.jp/ ZzzThai/Linux>, TE, Thai LaTeX tutorial, etc.

• By The group of students at The University of Electro-Communications, Tokyo. " Vuthichai’s Page " , http://www.ctrl.titech.ac.jp:80/˜ vuthi/ <http://www.ctrl.titech.ac.jp:80/

~vuthi/>

• Information about Thai computing by Mr. Vuthichai Ampornaramveth. " An annotated reference to the Thai implementations " , http://www.inet.co.th/cyberclub/trin/thairef/

<http://www.inet.co.th/cyberclub/trin/thairef/>

• Information about Thai character standard.

• By Mr.Trin Tantsetthi. " X window and Thai language " , http://members.xoom.com.cwg.x11th/ <http://members.xoom.com/cwg/

x11th/>

• By Mr.Rawat S.Pirom " SchoolNet Internet Sever " , http://www.school.net.th/linux-sis/ <http://www.school.net.th/

linux-sis/>

• Using Linux in School, Thailand.

• By NECTEC <http://www.nectec.or.th> (National Electronics and Computer Technology Center). " Thai Open Source Development " , http://members.xoom.com/inThai <http://members.xoom.com/

inThai>

• Mozilla Thai enabling.

• Open source Thai softwares and Libraries.

• By Mr.Samphan Raruenrom " Linux Thai Project " , http://www.geocities.com/SiliconValley/8302 <http://www.geocities.com/

SiliconValley/8302>

• Information about Linux in Thai language.

• By Kaiwal Development Team. " ThaiLinux unofficial Webboard " , http://lulu.mptc.eng.cmu.ac.th/HyperNews/get/ThaiLinux.html <http:

//lulu.mptc.eng.cmu.ac.th/HyperNews/get/ThaiLinux.html>

• Questions and answers about Linux in Thai language.

• By Mr.Pruet Boonma " Thai Linux installation project " , http://www.geocities.com/Tokyo/Bay/4521/ <http://www.geocities.

com/Tokyo/Bay/4521/>

• Installation guide in Thai language

4.3 FTP and Web sites

Most softwares and Thai fonts which introdued in this howto.

• ftp://ftp.fedu.uec.ac.jp/pub/thai/UEC/ZzzThai/Software/Linux <ftp://ftp.fedu.uec.ac.jp/pub/

     thai/UEC/ZzzThai/Software/Linux>
   • http://thaigate.nacsis.ac.jp/files/index.html <http://thaigate.nacsis.ac.jp/files/index.html>

• http://www.nectec.or.th/pub/software/i18n/thai <http://www.nectec.or.th/pub/software/i18n/

     thai>
Mule
   • ftp://etlport.etl.go.jp/pub/mule <ftp://etlport.etl.go.jp/pub/mule>
Ss
   • http://members.xoom.com/theera/ss/ <http://members.xoom.com/theera/ss/>
SunSite <ftp://sunsite.unc.edu/pub/linux/> and mirrors. doc/howto has the above mentioned HOW-

TOs. utils/nls and subdirectories contain files related to National Language Support. Developers should

take a look at locale-tutorial-0.8.txt.gz, locale-pack-0.8.tar.gz and cat-pack.tar.gz.

The GNU archives <ftp://prep.ai.mit.edu/pub/gnu/> has the recode package for character table conversion, the ABOUT-NLS file and the gettext package for locale support of some GNU applications and (of course) the latest versions of GNU emacs.

5 Acknowledgments and Copyright

Some parts of this HOWTO comes from The Linux Danish/International HOWTO by Thomas Petersen, petersen@post1.tele.dk (the original author) and Niels Kristian Bech Jensen, nkbj@image.dk.

Thank you to Phaisarn Techajaruwong for building Thai fonts and valuable discussion.

Thank you to Thai students at The University of Electro-Communications and Mr. Vuthichai Ampornaramveth for every help.

This HOWTO is copyrighted by Poonlap Veeratanabutr, poon-v@fedu.uec.ac.jp. It is distributed as other Linux HOWTOs under the terms described below.

Linux HOWTO documents may be reproduced and distributed in whole or in part, in any medium physical or electronic, as long as this copyright notice is retained on all copies. Commercial redistribution is allowed and encouraged; however, the authors would like to be notified of any such distributions.

All translations, derivative works, or aggregate works incorporating any Linux HOWTO documents must be covered under this copyright notice. That is, you may not produce a derivative work from a HOWTO and impose additional restrictions on its distribution. Exceptions to these rules may be granted under certain conditions; please contact the Linux HOWTO coordinator at the address given below.

If you have questions, please contact Tim Bynum, the Linux HOWTO coordinator, at

linux-howto@sunsite.unc.edu via email.